# A Non-Abelian Gauge Theory for Surface Excitations of ${}^{3}$e-B

### v1

@@ 4 Gauss Law and Superselection Rules
    Matter quantum numbers have been omitted in (4.16).
    Sticking to $SU(3)$, its Lie algebra has plenty of elements which do not commute with $Q_{3,8}\mathbb{I}$, for example $Q_{1}=Q(\mathbb{I}\lambda_{1})$ if $\lambda_{~{}I}\neq 0$. This will change the vector (4.16) to a new one, but preserving $c_{2,3}$ As no element of $\mathcal{A}_{loc}$ can change $\lambda_{3,8}\mathbb{I}$, this vector by GNS gives a new representation of $\mathcal{A}_{loc}$.
    The non-abelian symmetry $SU(3)_{L+R}$ is thus ‘spontaneously broken’ to $U(1)\times U(1)$ with generators $Q_{3,8}$ or to $U(2)$ if $I=0$ ( [4], [5], [6]).
@@
    A Remark on Effective Lagrangians and Anomaly Matching
@@
    Surface states in ${}^{3}$e-B
    Consider ${}^{3}$e $B$ in say a spherical container. ( Containers with other topologies such as toroidal can also be treated).
@@
    The field equations $F(\psi)=0$ involve only local fields $\psi$ and are gauge invaraint:
    $F(\psi)=0\Rightarrow F([Q(\Xi),\psi])=0$ (4.18)
    As $\Xi$ can be an $X$, (4.18) implies full gauge invariance even though $\psi$ may not commute with $Q(\Xi)$ as is the case with connections.
@@
    Adding to the $QED$ Hamiltonian $H_{0}$, we work with the new Hamiltonian $H=H_{0}+N$ which gives exactly the same equations of motion, but affects the partition functions.
    For such reasons , let us call $Q(\Xi)$ as chemical potentials. If $H_{0}$ is the ${}^{3}$e-B Hamiltonian we start with where the vector state is the ‘vacuum’
    $|0\rangle,\qquad Q(\Xi_{i})|0\rangle=0,$
    we can go to the sector with $H$ quantum numbers as in (4.16). This is the analogue of the excited states considered by Doplicher, Haag and Roberts [8]. Expectation values of local operators change on the folium of states (4.16) . An example is given below.
    Let $\mu$ be a constant with dimension of energy. We can then introduce a chemical potential in a new Hamiltonian
    $h=h_{0}+\mu Q(\Xi)$ (4.19)
    where $Q(\Xi)$ does not commute with the $Q_{3,8}$. Then the vector state (4.16) will change and so will the expectation of the generic surface operators $Q(\Xi^{\prime})$. Thus one has dynamics on the two sphere $S^{2}$.
    For example, suppose $Q(\Xi)=aQ_{1}+bQ_{5}$ with $a/b$ being irrational. The time $t=0$ inital state (4.16) changes with time and is not periodic since $\exp(i\mu Q(\Xi))$ has no period. One shows that it keeps evolving the diagonalised operators from one to another Cartan subalgebra [6] . ( One can prove that it never returns to the initial state of $t=0$. ).
    The spectrum of $h$ will keep changing with time. If $\mu$ is switched off at time $T$, there will be edge transitions to the ground state. May be they can be seen.
@@
    Interpretation of the Chemical Potential
    In $QED$, the term $N$ comes from coupling the system to an external source supplying the charge. The interpretation of $\mu Q(\Xi)$ is much the same, it comes from the measuring apparatus coupled to the system.
@@
    Expectation value of charge density $:J_{0}:$ vanishes in the vacuum sector, but it cannot vanish if $N$ is not an integral value $q$ in the charge $q$ sector. In this case, we have to construct $J_{0}$ from charged fields like the electron field. The last is not gauge invariant. It is associated with the field algebra.
    There are similar gauge invariant fields in any local gauge theory such as that of ${}^{3}$ He $B$. The operators $\mathrm{tr}(E^{2}(x)+B^{2}(x)),\mathrm{tr}E\cdot B$ are examples. Combination of such terms integrated gives the Hamiltonian in the absence of chemical potential and its expectation value keeps changing in time when the chemical potential keeps changing the vector state. Knowing their expectation values at time $t=0$ when the experiment starts and the chemical potential $\mu(t)Q(\Xi)$ is switched on ( it is now time independent ) and at time $T$, when it is switched off. one infers the properties of the local dynamics.
    An Important Remark
@@
    $S=\int d^{3}x~{}\mathrm{tr~{}}(-\frac{1}{4e^{2}}F_{\wedge}*F+A^{\mu}J_{\mu}).$ (4.20)
@@
    $\frac{1}{4e^{2}}F_{\wedge}*F+Q(A_{0}).$ (4.21)
    The standard treatments set $A_{0}=0$, but one sees that it plays the role of $\mu Q(\Xi)$. Its asymptotic value seems determined by experimentl setup: Having fixed the vector state, its evolution is governed by the chemical potential $Q(A_{0})$.
@@
    Likewise, the chemical potential $\mu Q(\Xi)$ should cause observable effects and help probing the system.
    5 Final Remarks
    This paper is restricted to considering surface effects. Still it suggests observable effects such as energy fluctuations and transitions due to them from emission of probably Goldstone modes . In addition, the approach predicts solitons [2]. It predicts as well theta vacua which are superselection sectors which can be constructed for ${}^{3}$e $B$ following [5] and which are spinorial. Such matters are being omitted for brevity.
@@ 5 Final Remarks
    6 Acknowledgements
    This paper could not have been written without the help of Arshad Momen. I thank him. Helpful comments from Arshad,l Parameswaran Nair,Al Stern and Rakesh Tibrawala are also appreciated. I also thank the Institute of Mathematical Sciences,Chennai and the hospitality of my colleagues there ( especially the Director Ravindran and S. Digal) where this work was completed.

### v4

@@ IV.1 Superselection Rules
Matter quantum numbers have been omitted in (27).
Sticking to $SU(3)$, its Lie algebra has plenty of elements which do not commute with $Q_{3,8}$, for example $Q_{1}=Q(\mathbb{I}\lambda_{1})$ if $\lambda_{~{}I}\neq 0$. This will change the vector (27) to a new one, but preserving $c_{2,3}$. As no element of $\mathcal{A}_{loc}$ can change $Q(\lambda_{3,8}\mathbb{I})$, this vector by GNS gives a new representation of $\mathcal{A}_{loc}$.
The nonabelian symmetry $SU(3)_{L+R}$ is thus ‘spontaneously broken’ to $U(1)\times U(1)$ with generators $Q_{3,8}$ or to $U(2)$ if $I=0$ ( [4], [5], [6]).
@@
IV.2 A Remark on Effective Lagrangians and Anomaly Matching
@@
IV.3 Surface states in 3He-B
Consider 3He $B$ in say in a spherical ball . ( Containers with other topologies such as solid torus can also be treated).
@@ IV.3 Surface states in 3He-B
The field equations $F(\psi)=0$ involve only local fields $\psi$ and are gauge invariant:
$F(\psi)=0\Rightarrow F([Q(\Xi),\psi])=0$ (29)
As $\Xi$ can be an $X$, (29) implies full gauge invariance even though $\psi$ may not commute with $Q(\Xi)$ as is the case with connections.
@@
Adding $N$ to the $QED$ Hamiltonian $H_{0}$, we work with the new Hamiltonian $H=H_{0}+N$ which gives exactly the same equations of motion, but affects the partition functions.
For such reasons , let us call $Q(\Xi)$ as generalised chemical potentials. If $h_{0}$ is the 3He-B Hamiltonian, we start with the ‘vacuum’
$|0\rangle,\qquad Q(\Xi_{i})|0\rangle=0,$ (30)
then we can go to the sector with $SU(3)_{L+R}$ quantum numbers as in (27). This is the analogue of the excited states considered by Doplicher, Haag and Roberts [9]. Expectation values of local operators change on the folium of states (27) . An example is given below.
Let $\mu$ be a constant with dimension of energy. We can then introduce a generalized chemical potential in a new Hamiltonian
$h=h_{0}+\mu Q(\Xi)$ (31)
where $Q(\Xi)$ does not commute with the $Q_{3,8}$. Then the vector state (27) will change in time and so will the expectation of the generic surface operators $Q(\Xi^{\prime})$. Thus one has dynamics on the boundary two sphere $S^{2}$.
For example, suppose $Q(\Xi)=aQ_{1}+bQ_{5}$ with $a/b$ being irrational. The time $t=0$ inital state (27) changes with time and is not periodic since $\exp(it\mu Q(\Xi))$ has no period. One shows that it keeps evolving the diagonalised operators from one to another Cartan subalgebra [6] .
One can prove that it never returns to the initial state of $t=0$. Its orbit is ergodic, being dense in the Grassmannian $Gr_{2}(C^{3})$
The spectrum of $h$ will keep changing with time. If $\mu$ is chosen to be time-dependent and is switched off at time $T$, there will be transitions to the ground state. Maybe they can be seen.
@@
IV.4 Interpretation of the Generalised Chemical Potential
In $QED$, the term $N$ comes from coupling the system to an external source supplying the charge. The interpretation of $\mu Q(\Xi)$ is much the same, it comes from the measuring apparatus coupled to the system. In a paper under preparation, this claim is confirmed using Takesaki duality and a measurement theory for guage fields is outlined.
@@ IV.4 Interpretation of the Generalised Chemical Potential
Expectation value of charge density $:J_{0}:$ vanishes in the vacuum sector, but it cannot identically vanish if $N$ , its integral, has the value $q$ in the charge $q$ sector. In this case, we have to construct $:J_{0}:$ from charged fields like the electron field. The last is not gauge invariant. It is associated with the field algebra.
There are similar gauge invariant fields in any local gauge theory such as that of 3He-$B$. The operators $\mathrm{tr}(E^{2}(x)+B^{2}(x)),\mathrm{tr}E\cdot B$ are examples. Combination of such terms integrated gives the Hamiltonian in the absence of chemical potential and its expectation value keeps changing in time when the chemical potential keeps changing the vector state. Knowing their expectation values at time $t=0$ when the experiment starts and the chemical potential $\mu(t)Q(\Xi)$ is switched on ( it is now time dependent ) and at time $T$, when it is switched off. one infers the properties of the local dynamics.
IV.5 An Important Remark
@@
$S=\int d^{4}x~{}\mathrm{tr}~{}(\frac{1}{2e^{2}}(\mathbf{E}^{2}-\mathbf{B}^{2})+A^{\mu}J_{\mu}).$ (32)
@@
$\int d^{3}x~{}\mathrm{tr~{}}\frac{1}{2e^{2}}(\mathbf{E}^{2}+\mathbf{B}^{2})+Q(A_{0}).$ (33)
The standard treatments set $A_{0}=0$, but one sees that it plays the role of $\mu Q(\Xi)$. Its asymptotic value seems determined by experimental setup. Having fixed the vector state, its evolution is also governed by the chemical potential $Q(A_{0})$.
@@
Likewise, the chemical potential $\mu Q(\Xi)$ should cause observable effects and help in probing the system.
V Final Remarks
This paper is restricted to considering surface effects. Still it suggests observable effects such as energy fluctuations and transitions due to them from emission of probably Goldstone modes.
In addition, the approach predicts solitons [2]. It predicts as well theta vacua which are superselection sectors which can be constructed for ${}^{3}$e-B following [5] and which are spinorial. Such matters are being omitted for brevity.
@@ V Final Remarks
Acknowledgements.
This paper could not have been written without the help of Arshad Momen and Babar Qureshi. I thank them. Helpful comments from Arshad, Parameswaran Nair, Al Stern and Rakesh Tibrewala are also appreciated. I also thank the Institute of Mathematical Sciences,Chennai and the hospitality of my colleagues there ( especially the Director Ravindran and S. Digal) where this work was completed.